OpenStudy (sidsiddhartha):

brain teaser: find the area of the region in which the complex number \[z(=x+iy) \] lies, under the following conditions \[z_1=\frac{ 2z-i }{ iz+2 }~~,|x_1|<1,y_1>0,|z_1|>1 \\\\\\where--\\z_1=x_1+iy_1\]

OpenStudy (anonymous):

So \(-1<x_1<1\) and \(y_1>0\)? No absolute value on the \(y\)?

OpenStudy (sidsiddhartha):

yeah that's all given no more :(

OpenStudy (ikram002p):

wanna work on this in morning :'(

OpenStudy (sidsiddhartha):

i was doing it like this-- \[z_1=\frac{ 2z-i }{ iz+2 } \\x_1+iy_1=\frac{ 2(x+iy)-i }{ i(x+y)+2 } \\=\frac{ 2x+i(2y-1) }{ (2-y)+ix }*\frac{ (2-y)-ix }{ (2-y)-ix } \\=\frac{ 2x(2-y)+x(2y-1) }{ (2-y)^2+x^2 }+\frac{ i[(2y-1)(2-y)-2x^2] }{ (2-y)^2 +x^2}\]

OpenStudy (sidsiddhartha):

@ikram002p :)

OpenStudy (anonymous):

Looks like \(z_1\) can be any number in this region.